The equilibrium concentrations of the reactants and products are [HA]=0.130 M , [H+]=3.00×10−4 M , and [A−]=3.00×10−4 M . Calculate the value of p
1
Expert's answer
2020-06-16T15:03:07-0400

"HA\to H^+ + A^-"

Equilibrium constant "=\frac{[A^-][H^+]}{[HA]}" "\frac{3\times 10^{-4}\times 3\times 10^{-4}}{0.130}=6.923\times 10^{-7}\ M"